OpenStudy (anonymous):

1.Create a quadratic function, f(x), in vertex form. The a should be between 4 and –4, the h will be your birth month, and the k will be your birth day. Write your equation below. 2.Using complete sentences, explain how to convert your birthday function into standard form. 3.Graph your function. Include your graph below. 4.Using complete sentences, explain how to find the average rate of change for f(x) from x = 4 to x = 7.

OpenStudy (anonymous):

my birthday is July twenty-ith (07/20)

OpenStudy (anonymous):

cool

OpenStudy (anonymous):

so................. do u still need help

OpenStudy (anonymous):

yes please

OpenStudy (campbell_st):

the vertex form is \[y = a(x - h)^2 + k\] you have k = 20 and h = 7 substitute them lastly choose a value for a remember its between -4 and 4 an easy choice is a = 1 a really each choice is a = 0... but your teacher probably won't like it.

OpenStudy (campbell_st):

then to get it to standard form... just distribute and get it into the form \[y = ax^2 + bx + c\] to graph it, use a table of values... given the vertex is at (7, 20) pick values either side so x : 3 4 5 6 7 8 9 10 substitute them to find y. then plot the points to get the parabola.